OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (5, -5) with 0° ≤ θ < 360°

OpenStudy (anonymous):

hey i actually got stuck @Astrophysics

OpenStudy (astrophysics):

Hey, no worries, where did you get stuck?

OpenStudy (astrophysics):

I will just put this for a reference \[x = r \cos \theta\]\[y= r \sin \theta\]\[r^2 = x^2+y^2 \implies r = \sqrt{x^2+y^2}\], and lets go over it

OpenStudy (anonymous):

ok so i plugged in everything... r=sqrt(5^2 + (-5)^2) but when i solved for theta my answer wasnt one of the options

OpenStudy (anonymous):

r = 5sqrt(2)

OpenStudy (astrophysics):

Ok lets see, so you did \[r = \sqrt{x^2+y^2} \implies r = \sqrt{(5)^2+(-5)^2} = \sqrt{50} = 5\sqrt{2}\] so far so good?

OpenStudy (anonymous):

yea i got that

OpenStudy (astrophysics):

Ok cool, so the angles can be quite tricky that's why I had drawn the triangle for you earlier, so here we have to use the tan ratio to find the angle

OpenStudy (anonymous):

ok so tan(theta) = -5/5

OpenStudy (anonymous):

???

OpenStudy (astrophysics):

Yes exactly!

OpenStudy (anonymous):

tan(theta) = -1 from here to I inverse it?

OpenStudy (astrophysics):

Yup \[\theta = \tan^{-1}(-1)\]

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

7pi/4

OpenStudy (astrophysics):

Good so that is 315 degrees

OpenStudy (anonymous):

ok so it would be 5sqrt(2),315

OpenStudy (astrophysics):

Yes, and the other you want for \[-5\sqrt{2}\]

OpenStudy (astrophysics):

Which is just 315-180

OpenStudy (anonymous):

135

OpenStudy (astrophysics):

Bingo

OpenStudy (anonymous):

so the answer is A?

OpenStudy (astrophysics):

Sounds good!

OpenStudy (anonymous):

Thank you so much for the help! :)

OpenStudy (astrophysics):

Np :)

OpenStudy (anonymous):

x is positive and y is negative so it is below x-axis x is positive , yis negative ,hence in fourth quadrant.

OpenStudy (astrophysics):

Yup I already had shown that in the previous post ;P

OpenStudy (anonymous):

you mind helping me with one more? sorry. lol

OpenStudy (astrophysics):

Sure haha

OpenStudy (anonymous):

Find all polar coordinates of point P = (2, 14°)

OpenStudy (astrophysics):

n is an integer

OpenStudy (anonymous):

x = 1.94059 y = 0.48384

OpenStudy (astrophysics):

Ok maybe I shouldn't have wrote that just ignore it and look at what I said after

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

so would it just be (2, 14 +/- 2npi) ???

OpenStudy (astrophysics):

Yes that works (2, 14 +360n) as we're using degrees

OpenStudy (anonymous):

is that it?

OpenStudy (astrophysics):

yes or (-2, 14+90+360n)

OpenStudy (anonymous):

wait im confused now where did the 90 come from?

OpenStudy (astrophysics):

Look at what I said twice, the polar representations

OpenStudy (anonymous):

so all the polar coordinates are (2, 14 +360n) , (-2, 14+90+360n)

OpenStudy (astrophysics):

Yes that sounds good

OpenStudy (anonymous):

ok thanks again. for all the help

OpenStudy (astrophysics):

The polar representation should be \[(r, \theta) = (r, \theta+2 \pi n)~~~\text{and}~~~(-r, \theta+(2n+1) \pi)\] I think I made a mistake, but your answers are right in any case

OpenStudy (anonymous):

oh ok. :)

OpenStudy (astrophysics):

wait wait that should be an OR not an AND haha ok

OpenStudy (anonymous):

lol its alright

OpenStudy (astrophysics):

xD

OpenStudy (astrophysics):

Ok have fun and take care haha

OpenStudy (anonymous):

you too. Thanks again

OpenStudy (astrophysics):

Np :)